OpenStudy (anonymous):

A line has a slope of −5 and contains the point (−2, 4). Which equations represent the line? Choose all answers that are correct. A. −5x + y = −6 B. 5x + y = −6 C. (y – 4) = −5(x + 2) D. (y + 4) = −5(x − 2)

OpenStudy (anonymous):

When given 2 points and a slope you can plug them into this equation y-y1 = m(x-x1)

OpenStudy (anonymous):

M is the slop, y1 is the y coordinate you have bin given and x1 is the x coordinate

OpenStudy (anonymous):

plug it in and you get y-(4)=-5(x-(-2))

OpenStudy (anonymous):

the negative in the x-x1 cancel out, so simplifying you will get y-4 = -5(x+2)

OpenStudy (anonymous):

Multiply out the -5 so you will have y-4=-5x - 10

OpenStudy (anonymous):

then move the negative 4 on the left side to the right to get y by itself and you will have y-4+4=-5x-10+4

OpenStudy (anonymous):

simplify and you have y=-5x-6

OpenStudy (anonymous):

now look at your options, none of them are in standard form but 2 of them look kind of similar

OpenStudy (anonymous):

we see the 6 is by itself and the y and x are together so rearrange your equation

OpenStudy (anonymous):

y+5x=-6 we just moved the x to the other side

OpenStudy (anonymous):

so the answer is b